# Compressing Chain-of-Thought in LLMs via Step Entropy

### v1

@@ Introduction
We first train the model on (problem, compressed CoT) pairs generated using our 80% entropy-based pruning strategy. The model learns to predict compressed reasoning paths and generate [SKIP] tokens by minimizing cross-entropy loss, providing robust initialization for reinforcement learning.
@@
While SFT teaches static imitation of compressed traces, it does not explicitly optimize the accuracy-efficiency trade-off. We employ Group Relative Policy Optimization (GRPO)(Shao et al. 2024) to further optimize this behavior through reward-driven learning.
@@
This two-stage process trains the model to strategically decide when to perform detailed reasoning versus when to skip steps, achieving efficient reasoning while preserving accuracy.
@@
ACC (%) Tokens ACC (%) Tokens ACC (%) Tokens ACC (%) Tokens
DeepSeek-R1-7B $78.54$ 393 497393\,497 $88.17$ 1 851 9171\,851\,917 $63.33$ 475 303475\,303 $35.71$ 546 097546\,097
SFT $78.47$ 223 767223\,767 $85.92$ 1 388 2411\,388\,241 $56.67$ 276 942276\,942 $30.00$ 353 154353\,154
SFT + GRPO $79.15$ 221 655 ($\downarrow$44%) $85.00$ 1 219 708 ($\downarrow$35%) $57.14$ 205 181 ($\downarrow$57%) $33.33$ 322 501 ($\downarrow$41%)
Table 3: Comparison of Pass@1 Accuracy (ACC %) and Thinking Tokens across baseline (DeepSeek-R1-7B), SFT, and SFT+GRPO training results on GSM8k, Math500, AIME2024, and AIME2025 benchmarks.
@@
We use DeepSeek-R1-Distill-Qwen-7B with 130k mathematical problems (DeepScaleR, OpenR1-Math) preprocessed using 80% entropy-based pruning, yielding 70k training samples after filtering sequences exceeding 4096 tokens. Stage 1 (SFT): 3 epochs on 70k samples using DeepSpeed Stage 2 and LoRA (r=16, $\alpha$=16). Stage 2 (GRPO): 10k samples with reward parameters $\tau_{high}=0.8$, $\tau_{low}=0.5$, $\tau_{skip}=100$, $\tau_{length}=3500$. Training uses DeepSpeed Stage 3, LoRA (r=16, $\alpha$=32), AdamW, G=14, KL=0.04 on 8×80GB GPUs. More details can be found in Appendix.
Experimental Analysis
@@
The results of our two-stage training process, presented in Table 3, demonstrate the effectiveness of our approach in creating an efficient yet powerful reasoning model.
The initial SFT stage successfully teaches the model a compressed reasoning style. Compared to the baseline DeepSeek-R1-7B model, the SFT model dramatically reduces the number of thinking tokens across all benchmarks. For example, on GSM8k, it cuts token usage by 43% while maintaining nearly identical accuracy. On AIME 2024, it achieves a 42% token reduction with a moderate drop in accuracy. This shows that the model effectively learns to use the [SKIP] token to bypass redundant steps identified in our entropy-based data preparation.
The GRPO stage further optimizes this behavior through reward-driven learning. The final two-stage training achieves impressive efficiency gains: 44% reduction on GSM8k, 35% on Math500, 57% on AIME 2024, and 41% on AIME 2025. Accuracy remains stable or improves slightly (GSM8k: 78.54%→79.15%), demonstrating that the model learns to skip truly redundant steps without compromising reasoning quality. The Math500 results reveal nuanced behavior where GRPO learns to be selectively less aggressive with compression to preserve accuracy, indicating adaptive reward-driven optimization rather than static imitation.
Overall, the results confirm the success of our training strategy, which provide a robust foundation for compressed reasoning and balance the trade-off between efficiency and task accuracy.
@@
ACC (%) Tokens ACC (%) Tokens
DeepSeek-R1-7B $78.54$ 390 545390\,545 $88.17$ 1 851 9171\,851\,917
$R_{c}+R_{sr}$ $75.93$ 734 913734\,913 $51.00$ 2 041 3512\,041\,351
$R_{c}+R_{sr}+R_{sn}$ $78.70$ 604 452604\,452 $85.00$ 1 346 9541\,346\,954
$R_{c}+R_{sr}+R_{sn}+R_{rl}$ $79.15$ 221 655221\,655 $85.00$ 1 219 7081\,219\,708
Reward Components Ablation Study
@@
Limitations
Despite the strong performance, our entropy-based CoT compression method has limitations. The 80% pruning threshold, while effective on DeepSeek-R1 and Qwen3 models, may not transfer to other architectures due to variations in reasoning redundancy. Furthermore, the method’s effectiveness is limited to mathematical reasoning tasks, necessitating validation for new application domains with potentially different optimal compression ratios. One possible way to address this challenge is to develop adaptive thresholds for task-aware and model-aware compression strategies, as discussed in extended experiments and discussions in Appendix.
Conclusion
We introduced a novel Chain-of-Thought compression framework that uses step entropy to identify redundant reasoning steps in LLM-generated CoTs. Our empirical validation demonstrates that pruning up to 80% of low-entropy steps maintains accuracy while achieving substantial token reductions (16-57% across benchmarks). Cross-model validation on DeepSeek-R1 and Qwen3 architectures confirms the broad applicability of step entropy as a generalizable principle for reasoning compression. Additionally, our two-stage training strategy enables models to autonomously generate compressed reasoning trajectories during inference. This work offers significant implications for efficient LLM deployment and provides new insights into the structure of reasoning processes.
Appendix A Appendix
Training Details
Stage 1: Supervised Fine-Tuning (SFT). The 70k pre-processed samples were used for SFT. This stage was conducted for 3 epochs using DeepSpeed Stage 2 for distributed training and LoRA PEFT (Hu et al. 2022) with parameters r=16 and $\alpha$=16.
Stage 2: Reinforcement Learning (RL). For the RL phase, a subset of 10k data samples was randomly selected from the 70k SFT-trained samples. We employed Group Relative Policy Optimization (GRPO) to further train the SFT-initialized model. This stage also utilized DeepSpeed Stage 3 for distributed training. The optimization objective involved a composite reward function designed to balance accuracy, [SKIP] token ratio, [SKIP] token number, and overall response length. LoRA PEFT was applied with parameters r=16 and $\alpha$=32, and the AdamW optimizer (Loshchilov and Hutter 2017) was used. Key GRPO parameters included G=14 samples per input and a KL coefficient of 0.04. All experiments were performed on a cluster of 8 GPUs, each has 80GB RAM and over 300 TFLOPS of BF16 compute performance.
Extended Experiments

### v2

@@ 3.3 Autonomous Compression via Two-Stage Training
We first train the model on (problem, compressed CoT, answer) pairs of the dataset in the Experiment 4.2 using our low entropy-based pruning strategy. The model learns to predict compressed reasoning paths and generate [SKIP] tokens by minimizing cross-entropy loss, providing robust initialization for reinforcement learning.
@@
While SFT teaches static imitation of compressed traces, it does not explicitly optimize the accuracy-efficiency trade-off. We employ Group Relative Policy Optimization (GRPO)(Shao et al., 2024) to further optimize this behavior through reward-driven learning.
@@
This two-stage process trains the model to strategically decide when to perform detailed reasoning versus when to skip steps, achieving efficient reasoning while preserving accuracy. The experiments of training and evaluation results are presented in Section 4.3, with an ablation study of the reward components available in Appendix F.
4 Experiments
This section presents experiments validating our entropy-based CoT compression method. We first conduct controlled studies to establish an optimal pruning ratio. We then demonstrate the effectiveness and generalizability of this strategy across multiple benchmarks and model sizes, justifying our step-level approach over token-level alternatives. Finally, we evaluate a two-stage training method that enables a model to autonomously generate compressed reasoning trajectories.
Table 1: Comparing the full COT baseline with our proposed step-entropy based pruning (Our) method, which prunes 80% of the lowest-entropy steps for DeepSeek-R1-7B, 14B and Qwen-8B. We conduct experiments to get the Pass@1 Accuracy(ACC)(%) and the number of Average Thinking Tokens per answer (contains the Unicode characters) during the inference on GSM8k, Math500, AIME2024 and AIME2025.
@@ 4 Experiments
ACC (%) Avg Tokens ACC (%) Avg Tokens ACC (%) Avg Tokens ACC (%) Avg Tokens
DeepSeek-R1-7B 78.54 298.33 88.17 3703.83 63.33 15843.43 35.71 18203.23
DeepSeek-R1-7B (Our) 80.82 294.29 (↓1.3%) 88.17 2604.23 (↓29.7%) 56.67 10092.80 (↓36.3%) 35.71 11471.17 (↓37.0%)
DeepSeek-R1-14B 82.64 283.63 84.37 2853.73 65.52 15414.83 58.62 18000.10
DeepSeek-R1-14B (Our) 84.00 278.29 (↓1.9%) 82.16 1980.97 (↓30.6%) 58.62 8705.57 (↓43.5%) 51.72 10842.07 (↓39.8%)
Qwen3-8B 94.46 3053.67 91.37 7138.49 79.31 20936.57 76.92 19902.43
Qwen3-8B (Our) 94.39 2557.47 (↓16.2%) 91.13 5209.24 (↓27.0%) 81.48 11533.57 (↓44.9%) 76.00 11716.63 (↓41.1%)
4.1 Determining the Optimal Pruning Ratio
To identify the safe threshold for step pruning, we conduct a controlled experiment using 50 samples from DeepScaleR (Luo et al., 2025) dataset on DeepSeek-R1-7B. And we investigate the impact on final answer accuracy by pruning steps using three distinct strategies with a mask ratio varying from 10% to 100% (“no-thinking” mode (Ma et al., 2025a)), shown in Figure 1 (left). In our methodology, pruned steps are replaced with a special “[SKIP]” token. This choice is informed by an ablation study detailed in Appendix E, which confirms that an explicit placeholder is more robust at high compression ratios than alternatives like direct removal.
  • Low-Entropy Steps Pruning: Steps with the lowest entropy are replaced by “[SKIP]”.
  • High-Entropy Steps Pruning: Steps with the highest entropy are replaced by “[SKIP]”.
  • Random Steps Pruning: Steps are replaced by “[SKIP]”. at random, serving as a control.
The results, illustrated in Figure 1, strongly validate our hypothesis. With Low-Entropy Steps Pruning, we observe that final answer accuracy remains stable and unaffected even when up to 80% of the lowest-entropy steps are masked. Beyond this 80% threshold, accuracy begins to decline, eventually converging to the accuracy of the “no-thinking” mode. This provides powerful evidence that a vast majority of low-entropy steps are indeed redundant. We found that with the best pruning strategy, we can prune up to 80% lowest-entropy steps (40% tokens redundancy) when not affecting the accuracy, shown in Figure 1 (right).
Conversely, with High-Entropy Steps Pruning, accuracy degrades immediately upon masking even a small fraction of steps. When the mask ratio exceeds 40%, performance drops below that of the “no-thinking” mode, indicating that removing these critical, high-information steps is more detrimental than providing no reasoning at all. The Random Steps Pruning strategy’s performance falls between the two, beginning to decline at a 40% ratio.
4.2 Validating Low-Entropy Steps Pruning Strategy
Models and Datasets.
We use models from the DeepSeek-R1 series (7B and 14B) and Qwen3-8B (Yang et al., 2025), which are open-source Large Reasoning Models with strong mathematical reasoning capabilities. To generate the initial CoT trajectories for our training data, we merged 40k samples from the DeepScaleR dataset (Luo et al., 2025) with 90k samples from the OpenR1-Math dataset (OpenR1-Math, 2025), creating a final composite dataset of 130k instances. To test the effectiveness and generalizability of our compression method, we evaluate performance on several standard mathematical benchmarks: GSM8k (Cobbe et al., 2021), Math500 (Lightman et al., 2023), AIME2024 (dataset card AIME, 2024) and AIME2025 (dataset card AIME, 2025).
Performance and Generalizability on Benchmarks.
Table 1 presents comprehensive results comparing our compressed CoT method against full CoT baselines. Our approach consistently achieves substantial efficiency gains while maintaining or improving accuracy across all models. The DeepSeek-R1 series shows remarkable token reductions: 29.7-37.0% for the 7B model and 30.6-43.5% for the 14B model across mathematical benchmarks, with GSM8k showing slight accuracy improvements for both sizes. Notably, Qwen3-8B demonstrates the strongest performance with impressive token reductions of 16.2-44.9% while maintaining competitive accuracy and even achieving slight improvements on AIME 2024 (79.31%→81.48%). The cross-architecture consistency—spanning both DeepSeek-R1 and Qwen3 model families—demonstrates that step entropy is a robust and generalizable principle for identifying redundancy, independent of model architecture, size. For a detailed analysis of our method’s performance across different domains, we conducted experiments on the MMLU benchmark (Hendrycks et al., ). The results, presented in Appendix G, demonstrate that step entropy is a generalizable metric for identifying redundancy beyond mathematical reasoning.
Scalability and Dataset Creation. To ensure our method scales beyond controlled experiments, we further validated that the 80% low-entropy pruning strategy holds at a much larger scale. To validate this strategy at scale, we applied this compression pipeline to the entire DeepScaler (40k) and OpenR1-Math datasets (90k). The results in Table 2 confirm that even across these tens of thousands of examples, the accuracy of the statically compressed CoTs remains almost identical to that of the full CoTs. This large-scale validation not only proves the robustness of our method but also serves as the direct procedure for Dataset Creation. The final dataset (130K), consisting of (problem, compressed CoT, answer) pairs our training strategy.
Model DeepScaleR OpenR1-Math
Full CoT 63.74 48.39
Compressed COT 62.17 47.31
Discussion of Our Method v.s. Directly Masking Tokens A crucial aspect of our methodology is the decision to prune entire reasoning steps rather than individual tokens. To justify this, we compared our step-based pruning approach against a token-based pruning baseline, where we remove the lowest-entropy tokens from the thinking trace irrespective of the steps they belong to. The results, shown in Figure 3, are unequivocal. While our step-pruning method maintains baseline accuracy even after removing up to 40% of the total thinking tokens, the token-pruning approach leads to a sharp and immediate decline in performance. Accuracy drops significantly after just a 20% token mask ratio. This demonstrates that a reasoning step is the correct semantic unit for compression. Removing individual low-entropy tokens (e.g., common words or operators) can break the syntactic and semantic integrity of a critical reasoning step, rendering it incomprehensible to the model. In contrast, removing an entire low-entropy step preserves the structure of the remaining, more important steps, leading to a much more robust compression strategy.
Figure 3: Comparing the accuracy of Our Method (via step-entropy) and Directly Masking Tokens (via token-entropy) across various thinking token mask ratios, with Full COT serving as the baseline of Deepseek-R1-14B on DeepScaleR dataset.
4.3 Training and Evaluation for Autonomous Compression
We evaluate our two-stage training framework (detailed in Section 3.3) for autonomous CoT compression. We analyze the trained model’s balance of accuracy and efficiency, and benchmark its performance against advanced compression baselines to demonstrate its state-of-the-art capabilities.
@@ 4.3 Training and Evaluation for Autonomous Compression
We use DeepSeek-R1-Distill-Qwen-7B with 130k mathematical problems (DeepScaleR, OpenR1-Math) preprocessed using 80% entropy-based pruning, yielding 70k training samples after filtering sequences exceeding 4096 tokens. Stage 1 (SFT): 3 epochs on 70k samples using DeepSpeed Stage 2 and LoRA (r=16, $\alpha$=16). Stage 2 (GRPO): 10k samples with reward parameters $\tau_{high}=0.8$, $\tau_{low}=0.5$, $\tau_{skip}=100$, $\tau_{length}=3500$. Training uses DeepSpeed Stage 3, LoRA (r=16, $\alpha$=32), AdamW, G=14, KL=0.04 on 8×80GB GPUs. More details and the descriptions of baselines can be found in Appendix C.
Training Experimental Analysis
@@
The results of our two-stage training process, presented in Table 3, demonstrate the effectiveness of our approach in creating an efficient yet powerful LRM.
[…]
Finally, comparing our trained model (SFT + GRPO) to the results of static pruning inference strategy from Table 1 reveals the key advantage of our training methodology. While static pruning is effective, our trained model learns a more dynamic and ultimately superior compression policy. On the most complex benchmark, AIME 2024, our trained model not only achieves a far greater token reduction (a 57.0% drop vs. 36.3% for static pruning) but also slightly improves accuracy (57.14% vs. 56.67%). This indicates that by training the model, it learns a more sophisticated, context-aware policy that can more effectively identify and prune redundancy in complex reasoning chains than a fixed, static rule. While static pruning preserves slightly higher accuracy on simpler benchmarks like Math500, the trained model’s ability to achieve state-of-the-art compression on challenging problems makes it a more robust and powerful solution for practical deployment.
Table 3: Comparison of Pass@1 Accuracy (ACC %) and Average Thinking Tokens per answer across baseline (DeepSeek-R1-7B), SFT, and SFT+GRPO training results on GSM8k, Math500, AIME2024, and AIME2025 benchmarks.
Method GSM8k Math500 AIME 2024 AIME 2025
ACC (%) Avg Tokens ACC (%) Avg Tokens ACC (%) Avg Tokens ACC (%) Avg Tokens
DeepSeek-R1-7B 78.54 298.33 88.17 3703.83 63.33 15843.43 35.71 18203.23
SFT 78.47 169.65 (↓43%) 85.92 2776.48 (↓25%) 56.67 9231.40 (↓42%) 30.00 11771.80 (↓35%)
SFT + GRPO 79.15 168.05 (↓44%) 85.00 2439.42 (↓35%) 57.14 6839.37 (↓57%) 33.33 10750.03 (↓41%)
Table 4: Comparative analysis of various Chain-of-Thought (CoT) compression methods on Math500 and AIME2024 benchmarks. All values represent the percentage change in accuracy (ACC) and generated token count relative to an uncompressed Full-CoT baseline.
Method Math500 AIME 2024
ACC Tokens ACC Tokens
Full-CoT (Baseline) - - - -
Token-efficient Prompts (Lee et al., 2025) $\downarrow$6.4% $\downarrow$11.1% $\downarrow$0.0 $\uparrow$7.2%
LC-Prompt (Xia et al., 2025) $\downarrow$1.6% $\downarrow$9.8% $\downarrow$2.6% $\uparrow$16.5%
CoT-Valve (Ma et al., 2025b) $\downarrow$10.6% $\downarrow$48.4% $\downarrow$15.0% $\downarrow$34.6%
TokenSkip (Xia et al., 2025) $\downarrow$5.2% $\downarrow$11.1% $\downarrow$12.3% $\downarrow$27.5%
R1-Compress (Wang et al., 2025a) $\downarrow$3.2% $\downarrow$20.3% $\downarrow$6.2% $\downarrow$12.9%
Our Method (SFT + RL) $\downarrow$3.2% $\downarrow$35.0% $\downarrow$6.2% $\downarrow$57.0%
Comparison with Baselines
To contextualize our performance, we compare our final trained model against several state-of-the-art CoT compression methods in Table 4. While most existing approaches offer some token savings, they often come at the cost of a significant drop in accuracy. For instance, CoT-Valve achieves a 48.4% token reduction on Math500 but with a steep 10.6% accuracy penalty. Our method, however, establishes a new state-of-the-art in the accuracy-efficiency trade-off. On Math500, our model matches the accuracy drop of the strongest baseline, R1-Compress (↓3.2%), while delivering nearly double the token savings (↓35.0% vs. ↓20.3%). The advantage is even more pronounced on AIME 2024. Our method again matches the accuracy performance of R1-Compress (↓6.2%) but achieves a massive 57.0% reduction in tokens—more than four times the savings offered by R1-Compress (↓12.9%). This robust performance across benchmarks validates that our entropy-guided, two-stage training strategy is highly effective at producing models that reason both accurately and efficiently, outperforming existing methods.
5 CONCLUSION and Future Work
Conclusion
We introduced a novel framework using step entropy to identify and compress redundant steps in LLM Chain-of-Thought reasoning. Our empirical validation, confirmed across multiple model architectures, demonstrates that pruning up to 80% of low-entropy steps maintains accuracy while achieving substantial token reductions (16-57%). Furthermore, our two-stage training strategy successfully enables models to autonomously generate these compressed reasoning trajectories, offering significant implications for efficient LLM deployment.
Limitations and Future Work
[…]
6 Acknowledgments
This work was supported in part by the CUHK Research Matching Scheme under Grant No. 7106937 and 8601130.
Ethics Statement
We have adhered to the ICLR Code of Ethics throughout the development of this research. Our work focuses on improving the inference efficiency of large language models on publicly available mathematical and multitask benchmarks (GSM8k, Math500, AIME, MMLU, etc.). We recognize the dual-use potential of highly efficient reasoning models; while they can serve as valuable tools for education and research, they could also be misused for tasks like automated academic dishonesty. The primary goal of our research is to enhance the practical deployment of LLMs by reducing their computational cost, making powerful reasoning tools more accessible and sustainable, and to contribute to the scientific understanding of redundancy in machine-generated thought processes. Our research utilizes publicly available pre-trained models. While our two-stage training process fine-tunes these models, we have made no effort to alter or remove their inherent safety mechanisms. We believe our work on creating more efficient and transparent reasoning systems aligns with the principles of responsible AI development.
Reproducibility Statement
We are committed to ensuring the reproducibility of our research. All experiments were conducted using publicly available large language models (DeepSeek-R1, Qwen3) and standard academic benchmarks, the specifics of which are detailed in Section 4 and the Appendix G. To facilitate the full reproduction of our findings, we will make our source code and training configurations publicly available upon publication. This release will include: (1) the implementation for calculating step entropy and performing static CoT pruning, (2) the scripts used to generate the compressed dataset for training, and (3) the complete code for both the Supervised Fine-Tuning (SFT) and Group Relative Policy Optimization (GRPO) training stages. Key hyperparameters and experimental settings, such as the pruning ratio ( $\kappa$=0.8), LoRA configurations, and the specific reward components for GRPO, are described in detail in our Experimental Setup of Section 4.3 and the Training Details in the Appendix C.
Appendix A The Use of Large Language Models
We use Large Language Models (LLMs), including ChatGPT and Gemini, solely for the purpose of editing and polishing the writing in this paper.
Appendix B Proofs of Theoretical Results
B.1 Proof of Lemma 1
Proof.
$I(S_{j};A|\bar{S}_{j})$ $=I(S_{j};A|S_{1},...,S_{j-1},S_{j+1},...,S_{N})$ (7)
$=H(S_{j}|\bar{S}_{j})-H(S_{j}|\bar{S}_{j},A)$ (8)
$\leq H(S_{j}|\bar{S}_{j})=H(S_{j}|S_{<j},S_{>j})$ (9)
$=H(S_{j}|S_{<j})-I(S_{j};S_{>j}|S_{<j})$ (10)
Since $I(S_{j};S_{>j}|S_{<j})\geq 0$, we have $I(S_{j};A|\bar{S}_{j})\leq H(S_{j}|S_{<j})$. ∎
B.2 Proof of Theorem 1
Proof.
Assume that there are K+1 steps, $\tilde{S}=S_{k_{0}},S_{k_{1}},...,S_{k_{K}}$, and we have
Without loss of generality, we assume the indices $k_{i}$ are arranged in descending order, i.e., $k_{i}<k_{i-1}$. Therefore, we could split the sequence $C/[S_{k_{0}},...,S_{k_{i}}]$ into $S_{<k_{i}}$ and $S_{>k_{i}}/[S_{k_{0}},...,S_{k_{i-1}}]$. Now consider the item with $k_{i}$, according to Lemma 1, we have:
$=H(S_{k_{i}}|S_{<k_{i}},(S_{>k_{i}}/[S_{k_{0}},...,S_{k_{i-1}}]))$ (13)
$\leq H(S_{k_{i}}|S_{<k_{i}})$ (14)
Therefore, we conclude that
$\leq\sum_{i=0}^{K}H(S_{k_{i}}|S_{<k_{i}})$ (16)
Appendix C Training Details
Stage 1: Supervised Fine-Tuning (SFT). The 70k pre-processed samples were used for SFT. This stage was conducted for 3 epochs using DeepSpeed Stage 2 for distributed training and LoRA PEFT Hu et al. (2022) with parameters r=16 and $\alpha$=16.
Stage 2: Reinforcement Learning (RL). For the RL phase, a subset of 10k data samples was randomly selected from the 70k SFT-trained samples. We employed Group Relative Policy Optimization (GRPO) to further train the SFT-initialized model. This stage also utilized DeepSpeed Stage 3 for distributed training. The optimization objective involved a composite reward function designed to balance accuracy, [SKIP] token ratio, [SKIP] token number, and overall response length. LoRA PEFT was applied with parameters r=16 and $\alpha$=32, and the AdamW optimizer Loshchilov and Hutter (2017) was used. Key GRPO parameters included G=14 samples per input and a KL coefficient of 0.04. All experiments were performed on a cluster of 8 GPUs, each has 80GB RAM and over 300 TFLOPS of BF16 compute performance.
Appendix D Baselines
To rigorously evaluate our proposed method, we compare it against several representative baselines spanning different Chain-of-Thought (CoT) compression strategies. These baselines include zero-shot prompting techniques that require no additional training, as well as more advanced, state-of-the-art methods that involve fine-tuning or specialized architectures.
  • Prompt-BeConcise: A zero-shot prompting technique, as explored in Lee et al. (2025), we select the prompt of Be Concise to encourage the model to generate a more succinct reasoning chain naturally.
  • LC-Prompt (Length-Control Prompt): Another zero-shot approach, adapted from Xia et al. (2025), that directly instructs the model to reduce the length of its reasoning by a fixed proportion, using the prompt of Please reduce about 50% of the words in the Chain-of-Thought process.
  • Advanced Compression Methods: We also compare against recent, more sophisticated techniques:
    CoT-Valve (Ma et al., 2025b): A method that fine-tunes a model to generate CoTs of a controllable length for adjusting reasoning verbosity.
    TokenSkip (Xia et al., 2025): An approach that enables controllable, token-level skipping during the generation process, allowing for fine-grained compression of the reasoning path.
    R1-Compress (Wang et al., 2025a): This method performs compression at a coarser granularity by operating on “chunks” of the reasoning chain, combining compression with a search algorithm to maintain logical coherence.
Appendix E Ablation Study on the Static Pruning Strategy
We conduct an ablation study to identify the optimal replacement strategy for pruned reasoning steps, as simply deleting them may disrupt the model’s logical flow. We compare four strategies at aggressive pruning ratios (80%, 85%, and 90%) on DeepSeek-R1-7B:
  • Replace with “[SKIP]”: Our proposed method, using a distinct special token as a placeholder.
  • Replace with “[MASK]”: A generic mask token.
  • Replace with “and then”: A natural language filler phrase.
  • Directly Remove Steps: Deleting the pruned step entirely from the context.
As shown in Figure 4, all methods maintain full CoT accuracy (74%) at an 80% pruning ratio. However, as compression increases, directly removing steps or using “[MASK]” leads to faster accuracy degradation (68% at 90% pruning) compared to using explicit placeholders like “[SKIP]” or “and then” (70% at 90% pruning).
This result indicates that explicitly signaling an omitted step is critical for reasoning coherence under high compression. We selected “[SKIP]” as our primary strategy due to its superior token efficiency (a single token) and its unambiguous semantic role as a placeholder, which avoids introducing potentially confounding natural language.
Figure 4: Ablation study on different replacement strategies for pruned low-entropy steps. The experiment is conducted on DeepSeek-R1-7B with the same sampled data of DeepScaleR in Fig 1 (left), comparing four methods for handling pruned steps at high pruning ratios (80%, 85%, 90%).
Appendix F Reward Components Ablation Study
@@
ACC (%) Avg Tokens ACC (%) Avg Tokens
DeepSeek-R1-7B $78.54$ $296.09$ $88.17$ $3703.83$
$R_{c}+R_{sr}$ $75.93$ $557.17$ $51.00$ $4082.70$
$R_{c}+R_{sr}+R_{sn}$ $78.70$ $458.27$ $85.00$ $2693.91$
$R_{c}+R_{sr}+R_{sn}+R_{rl}$ $79.15$ $168.05$ $85.00$ $2439.42$
@@
Appendix G Extended Experiments